OpenStudy (anonymous):

we are working a problem and we are to -8.620689 In (.91) We were wondering how do you put it into a TI-84 Plus?

OpenStudy (anonymous):

-8.620689*ln(.91)

OpenStudy (anonymous):

Hi! Good to see u again. I will try that real quick

OpenStudy (anonymous):

when I do that, I get .813023037

OpenStudy (anonymous):

the answer starts with 19

OpenStudy (anonymous):

If the answer starts with 19 I don't think that you have the right question.

OpenStudy (anonymous):

I will give you the question, so maybe that will make it easier

OpenStudy (anonymous):

t=-1/k Ln (1-c/m) and the formula is t=-1/k Ln (1-c/m)

OpenStudy (anonymous):

The question is: if mc's the theoretical maximum charge that a battery can hold and k is a positive that depends on the battery and the charger, the length of time + (in minutes) required to charge the battery to a given level c is given by....formula above

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

I don't see where those numbers came from though.

OpenStudy (anonymous):

it is the answer in the book: 19.8 we are practicing on the cal for an upcoming test.

OpenStudy (anonymous):

No no, I mean the numbers .91 and -8.620689

OpenStudy (anonymous):

oh. How would you work the problem?

OpenStudy (anonymous):

I don't know? I haven't seen what the problem is.

OpenStudy (anonymous):

the first problem is the "problem" t=-1/k Ln (1-c/m)

OpenStudy (anonymous):

we need to know how to enter on the cal a problem like -40 Ln (0.1) to get the answer 92.10340372

OpenStudy (anonymous):

ln .1 enter *-40 enter

OpenStudy (anonymous):

or just -40*ln(.1)

OpenStudy (anonymous):

I think the 84 has parens around the arguments of functions.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Yea! God bless